A race car can be slowed with a constant acceleration of -10 m/s2.
(a) If the car is going 43 m/s, how many meters will it travel before it stops?
1
Expert's answer
2014-03-15T03:08:29-0400
a=-10 m/s2
v0= 43 m/s
s0=0
s=s0 + v0*t+ a*(t*t)/2
a= (v(finish)-v0)/t-> t = -v0/a -> s=v0*t + a*(t*t)/2
= v0*(-v0/a) + a (v0/a)*(v0/a)/2
s= 92.45 m